OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! PLEASE HELP!!! Gino designed a garden for his rectangular backyard. The design he drew on paper is ten inches long. The length of his backyard is fifty feet. If his drawing is to scale and is 9 inches wide, how many feet wide is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard width. Show your work.

OpenStudy (anonymous):

@Hero @nincompoop @dan815

OpenStudy (anonymous):

@texaschic101 @iGreen

OpenStudy (anonymous):

12

OpenStudy (anonymous):

@edwards555 can you please explain how you got the answer? Like I had asked please?

OpenStudy (anonymous):

@jigglypuff314 @Michele_Laino

OpenStudy (anonymous):

@jigglypuff314 are you going to help me?

jigglypuff314 (jigglypuff314):

yeah, sorry I had been a little caught up in another problem I'm not all that good with words :/ so ummm it would kind of be like the ratio of a length to the other length is equal to the ratio of the width to the other width... and you would get 10/50 = 9/x solve for x and that would be your "yard width" :)

OpenStudy (anonymous):

I am sorry I am not sure I understand?

jigglypuff314 (jigglypuff314):

"the ratio of a length to the other length is equal to the ratio of the width to the other width" = (length) / (other length) = (width) / (other width) let (length) = 10 (other length) = 50 (width) = 9 (other width) = x and you will get 10/50 = 9/x from there you can solve for x and get your answer

OpenStudy (anonymous):

okay thank you, I have a few other questions if you are willing to answer and get more medals.

jigglypuff314 (jigglypuff314):

perhaps after I finish my own homework :P

jigglypuff314 (jigglypuff314):

I'm not here fore the medals tbh

OpenStudy (anonymous):

Oh okay well thank you for the help. I will post a new question for the next one and tag you there.